OpenStudy (anonymous):

plane flies along a straight line path after taking off, and it ends up 90.0 km farther east and 200.0 km farther north, relative to where it started. In what direction did it fly on the straight line path?

OpenStudy (math&ing001):

tan(x)=90/200 ie x=arctan(90/200)=24° So the plane took off 24° east.